Presented below are two independent situations.

(a) Sage Co. sold $1,890,000 of 12%, 10-year bonds at 105 on January 1, 2020. The bonds were dated January 1, 2020, and pay interest on July 1 and January 1. If Sage uses the straight-line method to amortize bond premium or discount, determine the amount of interest expense to be reported on July 1, 2020, and December 31, 2020. (Round answer to 0 decimal places, e.g. 38,548.)

Interest expense to be recorded $


(b) Pronghorn Inc. issued $570,000 of 9%, 10-year bonds on June 30, 2020, for $471,929. This price provided a yield of 12% on the bonds. Interest is payable semiannually on December 31 and June 30. If Pronghorn uses the effective-interest method, determine the amount of interest expense to record if financial statements are issued on October 31, 2020. (Round intermediate calculations to 6 decimal places, e.g. 1.251247 and final answer to 0 decimal places, e.g. 38,548.)

Interest expense to be recorded $
Answer #1

a) Premium on bonds payable = 1890000*5% = 94500

Interest paid = 1890000*12% = 226800

Premium amortization = 94500/10 = 9450

Interest expense to be recorded = 226800-9450 = 217350

b) Interest expense to be recorded = 471929*12%*4/12 = 18877
Answer #2

A.)

Issue = 1,890,000 * 1.05 = 1,984,500

Face value = 1,890,000

Premium of bond payable = 1,984,500 - 1,890,000 = 94,500

Premium amortized each period = 94,500 / (10 years x 2 payments) = 4,725


Interest payment = face value x interest rate 

= 1,890,000 x 0.06 = 113,400


Interest Expense = Interest payment - Premium amortization each period 

= 113,400 - 4,725 

= 108,675

B.)

Interest expense to be recorded:

= 471,929 * 0.12 * 4/12

=18,877


source: Correct Homework
answered by: anonymous